Give relation between current density and conductance?

Grade:12

1 Answers

Arun
6 years ago

According to Ohm's law, the potential difference between ends of a conductor is directly proportional to the current, that is

V α I

Or      V=IR                                             (2)

Where  R is resistance (that is obstruction offered to flow of changes)

Also, when a current I flows in a uniform cross-section area normally to it, then

I=∫J.ds

= ∫JdS cos 00

=J∫dS

I=JA          (3)     (where ∫dS = A)

Also, resistance                                            R= ρl/A                                                 (4)

Equation (3) can also be written as

R=l/σA                                             (5)

Where σ=1/ ρ is conductivity

Substituting equations (3) and (5) in equation (2) we get

V=JA*l/σA

V=Jl/σ

Or      J=σV/l                                       (6)

But electric field intensity E=V/l

By putting above value of E in equation (6),we get

J= σ E